OpenStudy (anonymous):

3/10 + 12/35 = ? In reduced terms

OpenStudy (anonymous):

you have to get both with the same denominator, so what will you do? :)

OpenStudy (anonymous):

find the LCD?

OpenStudy (anonymous):

could, but theres a much faster way

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

lol, be my guest elec ^_^

OpenStudy (anonymous):

without a calculator

OpenStudy (anonymous):

no need for a calculator lol

OpenStudy (anonymous):

would the answer be 4/5?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

it should be :\[\frac{225}{350} = \frac{45}{70} = \frac{9}{14}\] simplify ^_^

OpenStudy (anonymous):

so its 9/14?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

yes :) you multiply the first fraction with 35 and the other fraction with 10 ofcourse both the denominator and numerator to get the same denominator , then you simply add :) and start simplifying

OpenStudy (anonymous):

did that make any sense? ^_^"

OpenStudy (anonymous):

thank you so much =] im taking down all ur notes for the future =]

OpenStudy (anonymous):

^_^ good luck and no problem